### v2

# Training the Untrainable: Introducing Inductive Bias via Representational Alignment

Abstract
We demonstrate that architectures which traditionally are considered to be ill-suited for a task can be trained using inductive biases from another architecture. We call a network untrainable when it overfits, underfits, or converges to poor results even when tuning their hyperparameters. For example, fully connected networks overfit on object recognition while deep convolutional networks without residual connections underfit. The traditional answer is to change the architecture to impose some inductive bias, although the nature of that bias is unknown. We introduce guidance, where a guide network steers a target network using a neural distance function. The target minimizes its task loss plus a layerwise representational similarity against the frozen guide. If the guide is trained, this transfers over the architectural prior and knowledge of the guide to the target. If the guide is untrained, this transfers over only part of the architectural prior of the guide. We show that guidance prevents FCN overfitting on ImageNet, narrows the vanilla RNN–Transformer gap, boosts plain CNNs toward ResNet accuracy, and aids Transformers on RNN-favored tasks. We further identify that guidance-driven initialization alone can mitigate FCN overfitting. Our method provides a mathematical tool to investigate priors and architectures, and in the long term, could automate architecture design.
Project website at https://untrainable-networks.github.io
1 Introduction
When creating neural networks, as a community, we follow recipes that select among a few architectures that are known to work for particular tasks [63, 12, 25]. Architecture is critical, encoding essential inductive biases i.e. priors that profoundly impact their learning capabilities and performance across various tasks. Convolutional nets revolutionized vision [47, 31], and Transformers reshaped language [74, 20, 1]. Despite this, architecture design is a dark art because the precise relationship between architectures and the priors they impose is poorly understood. For example, there is discussion about exactly what the role of residual connections is [39]. This reflects a broader challenge: we rarely understand exactly what inductive biases our architectures encode. Our lack of understanding makes architecture design challenging. Given new application spaces for neural networks with rising compute costs like inference-time scaling [56], this challenge has become even more relevant.
Recent theorems [59] state that for each function which is efficiently Turing computable, there exists a deep network that can approximate it well. Furthermore, a graph representing such a function is compositionally sparse, that is the nodes of the associated Directed Acyclic Graph (DAG) represent constituent functions with a small effective dimensionality. A reasonable conjecture is that neural networks with an architecture which is similar to the DAG of the unknown target function are especially successful in learning it, as it is the case for convolutional networks for image recognition and similar tasks. However, empirically testing or transferring those structural priors remains an open challenge. Because we do not understand the relationship between the kinds of priors on the target functions that different architectures impose, even simple questions have no known answer. For example, can an FCN’s initialization be tailored to mimic a CNN’s inductive bias, despite their distinct graphs?
To bridge this gap, we introduce a novel empirical tool, guidance. Given a target network, we guide it with a guide network. In addition to the target’s original loss, the target attempts to match the representation of its intermediate layers to those of the guide. We use a measure of representational similarity [45, 17, 16], also termed a neural distance function, to compute the distance between representations of two arbitrary layers. Neural distance functions are often used in neuroscience to compare activity in networks and brains [68, 14, 71]. In light of recent work that shows that networks of very different architectures have internal activity that is extremely similar to one another [30, 14, 15], we repurpose this distance function as a means to transfer priors between networks layer by layer. Surprisingly, even a randomly initialized guide—incapable of solving the task—yields large performance gains, proving architectures alone encode powerful priors. This surprising finding demonstrates that neural architectures alone, independent of parameter training, impose meaningful inductive biases that are useful for downstream tasks.
Figure 1: Guidance makes untrainable networks trainable via representational similarity. Given a target which cannot be trained effectively on a task, we train this target with a layerwise representational‐alignment term against a fixed guide—trained or random—which remains unchanged during task training. This transfers only the guide’s architectural bias, turning a network that would otherwise overfit or underfit into one that learns effectively (e.g., a deep FCN guided by a random ResNet for image classification).
We make the following contributions:
  1. 1.
    We develop guidance to transfer priors between networks using representational alignment and investigate one representational alignment method, centered kernel alignment, CKA [45].
  2. 2.
    We empirically differentiate between architectural and trained inductive biases, showing that architectural priors alone can significantly improve network performance. This underscores the intrinsic structural power of architectures independent of learned parameters.
  3. 3.
    We show that RNNs significantly improve their copy-and-paste accuracy when guided by a Transformer. Transformers increase their parity accuracy when guided by an RNN. RNNs close much of the gap to a Transformer on language modeling when guided by one.
  4. 4.
    We show that deep or wide fully connected networks stop overfitting when guided by a ResNet. No-skip CNNs close much of the gap with ResNets when guided by a ResNet. Fully-connected networks stop overfitting with guidance-only initialization schemes.
Our method provides a powerful empirical tool to further theoretical insights in architectural design. Guidance enables systematic investigation of the structural foundations of successful architectures and clarifies the distinction between architectural and trained inductive biases. Minimizing CKA lets you clamp any subset of a target network’s activations onto those of a frozen guide network, sweep that clamp across architectures and priors, and observe what fails or succeeds. Guidance is a knob you can turn to inject architectural priors at will, something one cannot do with cross-entropy or weight-decay alone.
Our work has a number of limitations. We aimed for coverage of many tasks instead of maximal performance on any one task. This would have required us to carefully tune the hyperparameters involved. We preferred to show how guidance works in general rather than in cherry picked and carefully tuned settings. To that end, we also did not optimize networks to convergence, nor did we attempt to experiment with other optimizers. Once we reproduced a well-known problematic training phenomenon, we showed that it could be overcome. We consider a network trainable and a problem to be overcome when the original problem disappears. For example, successfully training fully connected networks for object recognition was hopeless because they immediately overfit; using our guidance method they no longer do so. This does not mean that they are necessarily useful as object recognizers at present. In the case of fully connected networks, their present performance with guidance training is too low, but with additional work we believe their performance could be substantially increased now that their train and test loss are moving in the right direction. In some cases, by applying guidance, we do see large useful improvements, such as with RNNs and Transformers, as well as deep CNNs, although much more remains to be exploited there too.
2 Related Work
Representational Distance: Our method builds on several metrics that measure distance between high-dimensional activations extracted from neural networks or activity in the brain [44]. Some of these distance metrics make comparisons based on kernel matrices [45, 17, 16] or relative distances [46, 55] between sample representations in a set. Others compute linear [75, 68] or orthogonal projections [5] from one set of representations to another. These metrics are designed based on a set of desired invariant properties such as permutation invariance or invariance to linear transformations.
Such approaches have been commonly applied in neuroscience for measuring representational distance of activations from networks and activity in the brain to understand which neural networks are architecturally most similar to the brain [75, 14, 71, 24]. Under this context, Han et al. [30] has shown the inability of current representational distance metrics – specifically the metric used here, centered kernel alignment – to distinguish representations based on architecture. This paper provides the foundation for our intuition that networks may have similar representations that allow for transferring inductive biases from one network to another.
Untrainable Networks: This work examines RNNs and transformers for sequence modeling and FCNs and plain CNNs for image classification. Prior work explored similar approaches but performed poorly compared to the guide networks we leverage for improved training.
In sequence modeling, classical RNNs [69, 58, 13, 29] were constrained by vanishing and exploding gradients [35], making them unsuitable for long sequence tasks requiring memorization [27]. Gradient flow techniques were developed, but significant progress came from architectures like LSTMs [34] and transformers [74]. Transformers, however, have been found untrainable on formal language tasks requiring full-sequence reasoning, where RNNs succeed [8].
For image classification, small feed-forward networks with 3-5 hidden layers and fewer than 100 units per layer were trained on object recognition datasets [53, 6, 41, 57]. These efforts prioritized training fit over generalization performance and achieved low results [53, 6]. Strategies to reduce overfitting, such as topological structure [67] or early stopping [9], were hindered by complex designs and hyperparameter tuning, leading to poor training fits. Further methods used alignment between thin deep FCNs and wide shallow FCNs to prevent overfitting, an approach similar to our paper [64]. Deep convolutional networks were also applied to image classification [47, 7] but struggled with vanishing gradients, limiting their depth.
Model Distillation: Guidance shares a resemblance with model distillation [33, 26, 66, 36]. Distillation transfers knowledge from a teacher model to a student model by introducing a new component to the loss function that enforces the student model to behave like the teacher model [42, 76]. This usually consists of penalizing the KL-divergence between the logit predictions of the student and teacher model.
Representation-based distillation [72, 10, 50] and alignment techniques have been proposed to improve alignment between two networks. Certain works have proposed correlation congruence or similarity preserving metrics [62] for aligning two networks, particularly as a way to do architecture search between CNNs [4]. Methods have been proposed that use CKA as an alignment approach between representations of two networks or with representations in the brain with notable improvement in network performance [65, 18].
We distinguish guidance from distillation. Guidance can use a smaller untrained guide instead of a larger trained teacher. This is due to guidance operating over intermediate activations of the network instead the output of the network probabilities or output features, like distillation does. Guidance also operates at many levels at the same time, aligning many layers at once. This helps address the credit assignment problem that gradient descent has when tuning weights early in a network. We also consider many more networks for guidance than is traditional for distillation including networks which have very different architectures like Transformers to RNNs. Distillation is usually carried out between two closely related architectures. We apply guidance to do the opposite.
3 Methods
Guidance introduces a term in the loss of a target network, $\mathcal{N}^{T}$, to encourage representational alignment with a guide network, $\mathcal{N}^{G}$. We update only the target’s parameters, $\theta^{T}$, while keeping the guide’s parameters, $\theta^{G}$, frozen. On each minibatch, we compute a similarity metric $\mathcal{M}$ (e.g., CKA) between guide‐layer activations $\mathbf{A}^{G}_{i^{G}}$ and matched target activations $\mathbf{A}^{T}_{i^{T}}$. We refer to the correspondence between layers of the guide $\{i^{G}\}$ and layers of the target $\{i^{T}\}$ as $I$. While this correspondence, $I$, could be complex as any two architectures can form a guide/target pair, here we choose architectures that make the correspondence obvious as is discussed later. For example, the stacked RNNs and Transformers have the same number of layers in our experiments.
The target and guide receive the same input. Per minibatch, we collect activations from intermediate layers of both networks. Layers of guide network are mapped to layers of the target network; see fig.˜1. We formulate the loss in terms of minimizing the representational dissimilarity, $\bar{\mathcal{M}}$, i.e., the complement of a representational similarity metric, between guide and target activations layer by layer, summing the results. Here we use linear CKA, though any differentiable similarity metric could plug into Eq. (1). Efficiency or incremental computation is much more important than it is in traditional applications since this operation happens for every minibatch. We discuss details in section˜B.1.
Given $\mathcal{L}_{T}$ as the original loss of the target network, the guide network’s original loss function is irrelevant. The guide could be pretrained on another task. In fact, it need not even have been trained at all, only its architecture shapes the target. This latter setting is what allows transferring architectural priors without transferring knowledge from the guide to the target, as there is none in a randomly initialized guide. See eq.˜1 for an overall loss. We discuss details in appendix˜B.
$\mathcal{L}(\theta^{T})=\mathcal{L}_{T}(\theta^{T})+\sum_{i\in I}\bar{\mathcal{M}}({\bm{A}}^{T}_{i^{T}}(\theta^{T}),{\bm{A}}^{G}_{i^{G}}(\theta^{G}))$ (1)
Equation˜1 minimizes a task loss while increasing alignment between the target and guide networks given the mapping between them. The mapping may be sparse; not every layer needs to be used. This is important for guidance with transformers or stacked RNNs, as will be explained later. Note that the guide’s parameters, $\theta^{G}$, are constants, i.e., the guide is never updated.
Metrics like CKA can capture and encode inductive biases in neural network computations. For instance, CKA is a measure that depends on second-order statistics, specifically pairwise sample distance matrices. Architectural choices imprint distinct features on those statistics. For example, consider local receptive fields in a convolutional layer. Units that cover neighboring pixels receive correlated input, and this is reflected in our activations. Such correlations are reflected in our distance matrices, and these can be transferred to distance matrices associated with FCN layers that lack local correlations. Similarly, weight sharing, where the same kernel is applied at every spatial location, will also be reflected in a distance matrix.
Layerwise Mapping
We design a simple method for mapping guide layers to target layers as part of providing supervision. The goal of this method is to make guide and target networks architecturally agnostic i.e. we can supervise any target network with any guide network.
As a simple approach, we evenly spread layer computations of our guide network over our target network. For example, if we consider ResNet-18 and a 50-layer FCN, we would map every convolutional ResNet layer to every second or third linear layer of the FCN. Intuitively, evenly spacing guide‐to‐target matches encourages the target to approximate the guide’s compositional functions. Through the design of evenly spreading layers of our ResNet-18, we are guiding the FCN to find a function similar to the guide network.
For our mapping, we consider activations from all tunable-weight layers (convolutional, linear, or RNN/LSTM). For multiple stacked RNNs, LSTMs, or transformers, we extract feature representations from intermediate layers in the stack as well. Using all layers is useful for guidance as it provides a strong signal to induce alignment between the target and guide networks during training. We empirically find that more layers leads to stronger results. Skipping layers based on non-linear transformations reduces memory overhead associated with storing representations per batch.
4 Experiments
We design several settings with different target and guide networks to thoroughly test our approach. We include a range of image and sequence modeling tasks. In choosing target networks, we consider a broad range of designs for networks that are not traditionally applied (e.g., a FCN in image classification).
To systematically evaluate our approach, we incorporate two settings. (1) Untrainable Architectures: Experiments where the target networks are difficult to train due to architectural limitations, irrespective of the task. For example, memory incorporation in RNNs or overfitting in deep FCNs. (2) Untrainable Tasks: Experiments where certain tasks are inherently challenging for specific architectures, making them untrainable without additional supervision. For example, sequence classification with transformers.
Tasks: We describe the task settings. We consider three sequence modeling tasks to allow for a broader range of architectural settings. We first consider a task called copy-paste [27]. In this task, we generate a sequence of numbers in the range of 1101-10. The model is trained to recover the same sequence in the output. In our setting, we consider sequence lengths that range from $20$ to $40$ values total (internal sequence and padding). We generate a copy-paste dataset, sampling sequences containing numbers between $1$ and $10$. We generate a total of $100,000$ examples, training on $80,000$ examples, validating on $10,000$ examples, and testing on $10,000$ examples.
We also include the parity task, a binary classification task where a model is fed a bitstring and outputs $1$ when there is an even number of ones in the bitstring and 0 otherwise. We generate a series of bitstrings with sequence lengths that range from $2$ to $50$ as done in prior work [8].
Finally, we consider a language modeling task using the WikiText-103 dataset [54] where models must predict the next token given some context. This uses the train, validation and testing splits defined by the WikiText dataset and for all experiments, we use a context length of $50$. We tokenize the text data using the GPT-2 [61] tokenizer.
For an image-based task, we focus on image classification and use the ImageNet-1K dataset [19] for training and testing. We use the splits defined by the dataset. We report accuracy on the validation set for all experiments.
Tasks Guide Networks Target Networks
Copy-Paste Transformer RNN
Parity RNN Transformer
Language Modeling (Small and Large) Transformer RNN
Image Classification ResNet-18 Deep FCN
Wide FCN
ResNet-50 Deep ConvNet
Table 1: Guide and target networks across tasks. Our network designs include several untrainable target networks and corresponding trainable guide networks.
Architectures: For all tasks, we describe our target untrainable architectures for each task separately as well as the guide networks that are employed to make the untrainable network trainable. We give an overview in table˜1. We provide further details in appendix˜C.
Sequence Modeling: For our copy-paste task, we use a vanilla, 4-layer RNN as our target network. In copy-paste, architectural and algorithmic limitations make RNNs an untrainable architecture. For our language modeling task, we include two settings with a small (4 layer) and large (6 layer, larger hidden dimension) RNN. In this setting, vanishing gradients and limited context incorporation make RNNs an untrainable architecture as the training loss saturates. For the parity task, we use a 1-layer transformer encoder architecture, similar to prior work [8, 28]. For the copy-paste task, we train a guide network, 4-layer transformer decoder model which achieves 96.90% accuracy. Similarly, for language modeling, we train a 4-layer transformer decoder guide network with a context window of 256. Our final perplexity is 34.15 for the small language modeling setting and 33.10 for the large language modeling setting. For the parity task, we train a 1-layer vanilla RNN as a guide network which achieves 100% accuracy as reported by [8].
Image Classification: We use three target networks: Deep FCN, Wide FCN, and Deep ConvNet. Deep FCN is a fully-connected network with 50 blocks consisting of feedforward layers followed by non-linearities. This network is an untrainable architecture, lacking inductive biases to prevent overfitting and having vanishing gradients. Wide FCN is a fully connected network with 3 blocks with feedforward layers that have 8192 units. This is categorized as an untrainable task due to a saturation in the training performance. Deep ConvNet is the same architecture as ResNet-50 [31], but without residual connections. This is categorized as an untrainable architecture due to the vanishing gradient problem. We use two guide networks: ResNet-18 and ResNet-50. ResNet-18/50 is a deep convolutional network with 18/50 convolutional blocks and residual connections. We refer to He et al. [31]. We supervise the Deep FCN and Wide FCN with ResNet-18 and supervise the Deep ConvNet with ResNet-50.
Training: For each setting, we train the base target network and perform an experiment where both a trained and untrained guide network supervises the base target network. All networks are trained with cross-entropy loss, without loss of generality. For all sequence modeling tasks, i.e. copy-paste, parity, and language modeling we use AdamW [52]. For language modeling, we also incorporate gradient clipping due to unstable training with long sequences. When training networks for image classification using ImageNet-1K, we use the Adam [43] optimizer.
To ensure consistency of comparisons across learning curves, we use a consistent batch size of 256. Representational similarity metrics are affected by the number of samples in the calculation, where more samples allows for the metric to approximate representational distance better. We use 256 as a proxy, dependent on GPU memory, although more memory would allow for bigger batch sizes with potentially better results. Due to the large number of training settings, we employ several different learning rates. We tune the learning rate carefully for baseline training to ensure maximal performance. We sweep the parameter across 5 different values and choose the results with the lowest validation loss. This ensures we are choosing the training with the best performance.
After choosing the optimal learning rate, we then train all networks and settings for 100 epochs with 5 random seeds to compute error bars. Our error bars are associated with the standard error across each step across all seeds. We choose the seed-based average test accuracy associated with the epoch with the lowest seed-based average validation loss.
5 Results
Experiment
Copy-Paste
Accuracy $(\uparrow)$
Parity
Accuracy $(\uparrow)$
Language Modeling
(Small) Perplexity()($\downarrow$)
Language Modeling
(Large) Perplexity $(\downarrow)$
RNN 14.35 $\pm$ 0.01 100 69.19 $\pm$ 1.89 89.13 $\pm$ 2.00
Untrained RNN 2.32 $\pm$ 0.41
Transformer 96.98 71.98 $\pm$ 3.16 34.15 33.10
Untrained Transformer 1.04 $\pm$ 0.81 5.19e5 $\pm$ 90.44 5.19e5 $\pm$ 90.44
RNN $\rightarrow$ Transformer 78.49 $\pm$ 2.16
Untrained RNN $\rightarrow$ Transformer 70.38 $\pm$ 4.17
Transformer $\rightarrow$ RNN 23.27 $\pm$ 1.02 40.01 $\pm$ 1.54 36.91 $\pm$ 1.04
Untrained Transformer $\rightarrow$ RNN 42.56 $\pm$ 1.51 59.61 $\pm$ 2.33 47.17 $\pm$ 2.50
Table 2: Guidance improves performance for sequence modeling. RNN performance improves dramatically when aligning with the representations of a Transformer for copy and paste, as well as for language modeling with small and large RNN architectures. RNNs close most of the gap to Transformers for language modeling and are likely competitive with further scale. Transformers in turn, improve parity performance when aligning with an RNN. Guidance is able to transfer priors between networks.
Experiment ImageNet Top-5 Validation Accuracy $(\uparrow)$
ResNet-18 89.24
Untrained ResNet-18 0.24 $\pm$ 0.043
ResNet-50 92.99
Untrained ResNet-50 0.54 $\pm$ 0.029
Deep FCN 1.65 $\pm$ 1.21
ResNet-18 $\rightarrow$ Deep FCN 7.50 $\pm$ 1.51
Untrained ResNet-18 $\rightarrow$ Deep FCN 13.10 $\pm$ 0.72
Wide FCN 34.09 $\pm$ 0.91
ResNet-18 $\rightarrow$ Wide FCN 43.01 $\pm$ 0.92
Untrained ResNet-18 $\rightarrow$ Wide FCN 39.47 $\pm$ 0.31
Deep ConvNet 70.02 $\pm$ 1.52
ResNet-50 $\rightarrow$ Deep ConvNet 78.91 $\pm$ 2.16
Untrained ResNet-50 $\rightarrow$ Deep ConvNet 68.17 $\pm$ 2.54
Table 3: Guidance improves performance for image classification. Alignment with a ResNet dramatically improves a deep FCN, particularly with an untrained ResNet. Significant gains are seen with a wide FCN as well. Deep CNNs without residuals gain only with a trained ResNet. Across all settings, guidance can help train architectures that were otherwise considered unsuitable.
Sequence Modeling: On the copy-paste task, guiding a 4-layer RNN with a Transformer improves copy-paste accuracy by over 25%. See fig.˜2 and table˜2 Previous studies blamed RNN failures on vanishing gradients and memorization limits. Our results show a potential optimization scheme for RNNs that is applicable for sequence memorization. Remarkably, a random (untrained) Transformer guide outperforms a trained one, suggesting pure architectural bias drives gains. We believe this is because optimization with randomly initialized networks is easier due to the degrees of freedom in CKA. See appendix˜E. We plan to explore this more thoroughly in further analyses. These gains persist under our layerwise ablations and metrics; see appendix˜K and appendix˜G.
On the parity task, a 1-layer Transformer guided by an RNN improves its test accuracy by 7%. This is a complementary result to copy-paste and language modeling where the guide network was a transformer and our target network was a RNN. This improves over results from several prior papers [8] that have pointed out fundamental limitations of transformers to perform formal language tasks.
Unlike copy-paste, the performance improves when using a trained RNN as the guide network. This could be due to the wide gap in performance between an untrained RNN and trained RNN on parity. Parity uniquely benefits from learned positional encodings in the trained RNN, which the Transformer lacks. This information is likely crucial to the transformer, which has limited sequence pooling capacity and fewer degrees of freedom.
On language modeling, similar to copy-paste, guided RNNs halve the perplexity from ~70 points to 35 points on WikiText-103, closing in on Transformer baselines. While performance generally saturates for the 4-layer RNN, guidance continuously improves the RNN performance by over 30 points for text perplexity for both trained and randomly initialized guide networks. Scaling up to a 6-layer RNN further cuts perplexity by 10 points, indicating guidance scales with model size. This implies that information from the transformer can be transferred to the RNN. We also believe that this has exciting implications for scaling laws with RNNs. We see a similar trend with a randomly initialized transformer as the guide network, implying that architectural priors in the transformer are driving improvement in guided network performance.
Figure 2: Training and validation under guidance for all experiments reported in table˜1. For every result in Table˜3 and Table˜2, we show the training and validation loss with error bars across multiple runs, although these are often too small to see. Note that often the best results occur with the untrained guide.
Image Classification: Guidance boosts validation accuracy by 5–10% across our Deep FCN, Wide FCN, and Deep ConvNet; see table˜3. We also observe significantly better loss curves from a better fit with the training loss and reduced overfitting with the validation loss. An untrained ResNet-18 guide outperforms its trained counterpart on Deep FCN, underscoring pure architectural priors like with copy-paste. For example, the Deep FCN results in the top left of fig.˜2 are significantly better with a randomly initialized ResNet-18 as the guide network instead of a trained ResNet-18. This trend also occurs with Wide FCN. We show results with other neural distance functions in appendix˜J.
The Deep ConvNet (no skips) gains only from a trained ResNet-50 guide—implying residual connections require learned weights to shape representation. This explanation provides an additional interpretation for the role of residual connections and their influence on the representation space. This indicates that residual connections must be trained to have an influence on the representation space. This aligns with prior studies of residual connections [39, 32].
Figure 3: Guidance aligns error consistency. The relationship between the guide networks is mirrored in that of the guided networks, even when the target is entirely unlike the guides initially. This is additional evidence that guidance doesn’t just improve performance arbitrarily; the target becomes more like the guide.
Error Consistency: Guided FCNs mirror the ResNet–ViT error overlap, proving they inherit the guide’s decision patterns. Using Deep FCN as our target model, we guide it with a ResNet-18 or a ViT-B-16 [21]. We then measure the error consistency [23] between all of the networks; see fig.˜3. The error consistency between the initial FCNs is entirely unlike the ResNet-18 or ViT-B. Guidance creates two FCNs which have the same relationship to one another. The ResNet-18-guided FCN and ViT-B-guided FCN have the same error consistency with respect to one another as ResNet-18 and ViT-B do. It’s not just that the FCN gets generically better; it adopts a prior from the original architecture. We provide further details of the error consistency metric in appendix˜F.
Figure 4: Initializing fully connected networks with guidance can overcome overfitting. First, we align a Deep FCN to a random ResNet-18 on noise for 300 steps, then train normally. This two-stage scheme mirrors full guidance, and leads to a similar performance improvement. This suggests that FCNs have guidance-inspired initializations that avoid overfitting.
Network Initialization: Is guidance needed throughout training, or is the effect of the guide to move the target into a regime where the two are aligned and the target can be optimized further without reference to the guide? The answer to this question can shed light on whether the guide is telling us that better initializations are likely to exist for the target. To answer this question, we minimize the representational dissimilarity between our target and guide network for a nominal number of training steps, 300. Then we apply task training on the resulting target network with no guidance. Pre-aligning FCN layers to a random guide for 300 steps stops overfitting entirely—no ongoing guidance needed; see fig.˜4. Furthermore, while preventing overfitting, we have lower training loss from guidance, indicating a better fit. This implies that there exists a better initialization for FCNs.
Figure 5: Guidance outperforms distillation: We include a comparison between guidance and distillation for all settings with trained and untrained guide networks/teacher networks. We find that guidance outperforms distillation in all settings, highlighting that, unlike guidance, distillation fails in settings with an untrained teacher.
Guidance vs Distillation: We include a comparison between guidance and knowledge distillation [33] in fig.˜5. We find that guidance improves significantly over distillation, particularly when the teacher network is untrained. This is significant, demonstrating that guidance can exploit untrained networks for transferring inductive biases. This indicates that matching internal representations provides a much stronger signal over just matching output behavior. We discuss further in appendix˜H.
6 Conclusion
We demonstrated that guidance eliminates the failure modes of networks previously considered unsuitable or ineffective for specific tasks. Aligning with another network overcomes these shortcomings by transferring inductive biases—either architectural and knowledge-based, or solely architectural when using an untrained guide. This allows guidance to distinguish tasks and architectures that are dependent on architectural biases rather than learned biases. We provide further explanations and intuition for guidance in appendix˜I. There are many potential aspects that may distinguish between the effects of architectural and inductive biases in improving performance of these architectures, which we aim to explore in future work.
This also opens the door to many applications. Our method can be used to study representational and functional design of neural networks in new ways to reanalyze prior theory of neural network optimization. For example, we can understand distances between architectural components based on which target networks are easier to guide with a particular guide network. We also refine this notion to include a narrow channel through which guidance can occur, the representational similarity. This can serve as a kind of probe.
We believe guidance also reveals a new conceptual lens on inductive biases and architectures. When we use an untrained guide network, we reveal what an architecture by itself brings to the table. Guidance with a trained guide reveals how much learned representations can change performance. These are different phenomena whose root causes are not understood at present but that could elucidate the relationship between priors and architectures.
Our results suggest practical applications by significantly narrowing the performance gap between vanilla stacked RNNs and Transformers and showing signs of scaling, albeit in small-scale experiments with 150M parameters or fewer. Given that stacked RNNs are equivalent to single-layer RNNs, most directly to delayed RNNs [73], this implies that complex modifications to RNNs may be unnecessary for language modeling. For other networks, like fully connected ones, we only overcame the initial obstacle. Further research is needed to refine these into effective vision models, as they avoid immediate overfitting. In the future, we hope to find methods for making these networks competitive with their guide networks.
Guidance also proved to be a tool with which to discover the possibility of new initializations. At the moment, no known method exists to find better initializations for networks. In some cases as with the FCNs for vision, guidance can be disconnected after a nominal number of steps, but still goes on to regularize the target network. This strongly implies that an initialization regime for that target with the same regularization exists. This is all that guidance could do in that case. We now need tools to go backwards, given networks which are correctly initialized and networks which are not, discover what that initialization is. This is a much better place to be in. A systematic sweep of targets and guides to look for better initializations should be carried out.
Looking into the long-term future, guidance invites us to treat architecture itself as a trainable prior added directly to a generic network’s loss. Because guidance can rescue models that previously overfit or underfit, we can revisit designs abandoned during neural architecture search. These threads demonstrate the major possibility for future work in this space. Guidance is a tool, not a finished and well-understood theory or doctrine. Tools are useful for enabling more discoveries. We believe that these discoveries will allow the community to more easily pursue questions that training difficulties might have prevented in the past, especially with regard to understanding the relationships between architectures.
Acknowledgements
This work was supported by the Center for Brains, Minds, and Machines, NSF STC award CCF-1231216, the Brains, Minds, and Machines Summer School, the NSF award 2124052, the MIT CSAIL Machine Learning Applications Initiative, the MIT-IBM Watson AI Lab, the CBMM-Siemens Graduate Fellowship, the DARPA Mathematics for the DIscovery of ALgorithms and Architectures (DIAL) program, the DARPA Knowledge Management at Scale and Speed (KMASS) program, the DARPA Machine Common Sense (MCS) program, the Department of the Air Force Artificial Intelligence Accelerator under Cooperative Agreement Number FA8750-19-2-1000, and the Air Force Office of Scientific Research (AFOSR) under award number FA9550-21-1-0399. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Department of the Air Force or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein. V.S. and D.M. are supported by the National Science Foundation Graduate Research Fellowship under Grant No. 2141064. Any opinion, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Science Foundation.
Appendix A Appendix Overview
We present additional details of guidance, experiments and analysis, as well as additional results. In Appendix˜B, we provide additional details for our guidance approach, with a full explanation of centered kernel alignment. In Appendix˜C, we provide additional details on our architectural designs and network training. In Appendix˜D, we introduce a new experiment where we feed noise to our guide network rather than real data and compute representational alignment, leading to similarly improved results. This further establishes a transfer of a prior rather than knowledge. In Appendix˜E, we show visualizations of the representational dissimilarity loss over training to give context of dynamics over training and show additional explanations for results with trained and randomly initialized guides. In Appendix˜F, we provide further explanation of error consistency as a measure of functional similarity between networks. In Appendix˜G, we provide test accuracy metrics over training as a complementary of network performance over training outside of cross-entropy loss. In Appendix˜H, we provide a comparative baseline to guidance, distillation [33]. We find that basic distillation performs worse in comparison. In Appendix˜I, we provide an interpretation and analysis of guidance to better characterize and understand results based on changes in the internal geometry of a network before and after guidance. In Appendix˜J, we apply guidance with additional neural distance functions including RSA [46] and ridge regression. We find a correlation between the success of our approach and degrees of freedom of a particular distance function. In Appendix˜K, we provide additional ablations over guidance such as guiding a certain number of layers or guiding specific layers of a network.
Appendix B Methods Overview
We give an overview of guidance in algorithm˜1 and highlight crucial changes to base neural network training in either red or blue. We use blue to indicate the collection of network activations and red to indicate the layerwise mapping and representational alignment using a distance metric. This gives an overview of our layer mapping between the target and guide network. Crucially, we find that the simplest layer mapping where we evenly distribute guide network layers across target network layers for supervision obtains strong results.
Algorithm 1 Guidance: Guide Network Representational Alignment
1:Target network; $\mathcal{N}^{T}$ with parameters $\theta^{T}$; Guide network $\mathcal{N}^{G}$; Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{j=1}^{N}$; Representational Distance Metric $\bar{\mathcal{M}}$; Loss function $\mathcal{L}^{T}$
2:for $j=1\to N$ do
3:  # Base training with vanilla loss function
4:  $\text{outputs}\leftarrow\mathcal{N}^{T}(x_{j})$
5:  $\text{loss}\leftarrow\mathcal{L}_{T}(\text{outputs},y_{j}~|~\theta^{T})$
6:  # collect layer activations
7:  $\{{\bm{A}}_{i^{T}}^{T}\}_{i^{T}=1}^{t}\leftarrow\text{activations}(\mathcal{N}^{T}(x_{j}))$
8:  $\{{\bm{A}}_{i^{G}}^{G}\}_{{i^{G}}=1}^{l}\leftarrow\text{activations}(\mathcal{N}^{G}(x_{j}))$
9:  # Get step size between the number of layers between the two networks for layer mapping.
10:  if $l>1$ then
11:    $step\leftarrow(t-1)/(l-1)$
12:  else
13:    $step\leftarrow 1$
14:  end if
15:  # Map the layers and add up layer-wise representational distance
16:  $\text{total}\leftarrow 0$
17:  for $i=1\to l$ do
18:    $\text{index}\leftarrow\min(\text{round}(i\times step),t-1)$
19:    $\text{rep}\leftarrow\mathcal{M}({\bm{A}}_{\text{index}}^{T},{\bm{A}}_{\text{index}}^{G})$
20:    $\text{total}\leftarrow\text{total}+\text{rep}$
21:  end for
22:  $\text{loss}\leftarrow\text{loss}+\text{total}$
23:end for
B.1 Centered Kernel Alignment
To compare representations, we use a representation similarity metric, $\mathcal{M}$, which corresponds to centered kernel alignment (CKA) [45, 16, 17] in our setting. We specifically consider linear CKA.
CKA uses kernel functions on mean-centered representations to compute representational similarity matrices, which are then compared via the Hilbert-Schmidt Independence Criterion (HSIC). More specifically, suppose we have two sets of representations ${\bm{R}}\in\mathbb{R}^{b\times d_{1}}$ and ${\bm{R}}^{\prime}\in\mathbb{R}^{b\times d_{2}}$. We first compute the Gram matrices for each set of representations
${\bm{K}}={\bm{R}}{\bm{R}}^{T},{\bm{L}}={\bm{R}}^{\prime}{\bm{R}}^{\prime T}$ (2)
We center the Gram matrices by introducing a matrix, $H$, where $H={\bm{I}}_{b}-\frac{1}{n}{\bm{1}}{\bm{1}}^{T}$.
$\tilde{{\bm{K}}}={\bm{H}}{\bm{K}}{\bm{H}},\tilde{{\bm{L}}}={\bm{H}}{\bm{L}}{\bm{H}}$ (3)
We compute the HSIC on the Gram matrices.
[…]
Finally, we define our linear CKA metric as:
$\mathcal{M}({\bm{R}},{\bm{R}}^{\prime}):=\text{CKA}({\bm{K}},{\bm{L}})=\frac{HSIC({\bm{K}},{\bm{L}})}{\sqrt{HSIC({\bm{K}},{\bm{K}})*HSIC({\bm{L}},{\bm{L}})}}$ (5)
In our setting, we consider representational dissimilarity and aim to minimize the dissimilarity between representations from our target network and guide network. We define this as:
$\bar{\mathcal{M}}({\bm{R}},{\bm{R}}^{\prime})=1-\mathcal{M}({\bm{R}},{\bm{R}}^{\prime})$ (6)
Linear CKA ranges from 0 (identical representations) to $1$ (very different representations). Because of this, we take the complement by subtracting the linear CKA from $1$ to represent dissimilarity.
B.2 Methodology Limitations
Our guide network supervision through representational alignment has one primary limitation due to increased memory usage during training. Due to saving activations across several layers of the two networks, GPU memory usage increases dramatically. Moreover, our methodology works better as batch size increases since this allows for better approximation of representational similarity, increasing memory usage even more. Furthermore, including more layers for supervision leads to improved results.
In this paper, we introduce simple techniques to handle memory constraints such as gradient accumulation and gradient checkpointing [51]. In practice, more memory optimization techniques may become necessary to consider larger untrainable networks. Further work could consider using stronger representational alignment strategies to reduce the number of samples necessary to achieve a strong fit.
Appendix C Architecture and Training Details
C.1 Architectural Design Details
For all tasks, we describe our target untrainable architectural designs for each task separately as well as the guide networks that are employed to make the untrainable network trainable.
C.1.1 Copy-Paste
Target Networks
RNN: We design a 4-layer RNN with a hidden dimension of 768 units, followed by a fully connected layer. In copy-paste, architectural and algorithmic limitations make RNNs an untrainable architecture for the task. Specifically, RNNs must memorize the input sequence which is difficult, particularly with a padding token. RNNs are generally considered to be inapplicable to the copy-paste task.
Guide Networks
Transformer: We consider a 4 layer transformer decoder architecture with a hidden dimension of 768 units across 12 transformer heads. The transformer is well-suited for copy-paste as the attention mechanism can act as a routing mechanism for the sequence. We train the transformer guide from scratch, as with language modeling and achieve $96.90\%$ accuracy on the task.
C.1.2 Parity
Target Networks
Transformer: Similar to Bhattamishra et al. [8], we design a 1 layer transformer encoder network with a hidden dimension of 64 units across 4 attention heads. Transformers have lower accuracy on formal language tasks that require reasoning over a sequence in comparison to traditional sequence models [28]. Due to the enormous gap in performance and saturation of performance, we categorize the transformer as untrainable.
Guide Networks
RNN: We include a 1 layer vanilla RNN with a hidden dimension of 64 units. Similar to Bhattamishra et al. [8], we achieve 100% accuracy on the task.
C.1.3 Language Modeling
We include two language model settings to test scaling in RNNs. The first setting uses small networks, which we refer to as Small RNN and small Transformer. The other uses a large RNN and large Transformer.
Target Networks
Small RNN: We design a 4 layer vanilla RNN with a hidden dimension of 512 and with a ReLU activation function. We train this on sequences with a context length of 75. This makes the network untrainable due to problems associated with exploding and vanishing gradients during backpropagation through time.
Large RNN: We design a 6 layer vanilla RNN with a hidden dimension of 1024 and with a ReLU activation function. We train this on sequences with a context length of 128. Prior work has demonstrated that larger RNNs are difficult to train in practice [49].
Guide Networks:
Small Transformer: We design a 4 layer transformer decoder network with 16 attention heads and a hidden dimension of 512. We train the transformer on WikiText-103 with a context length of 256 and achieve a final test perplexity of 34.15.
Large Transformer: We design a 4 layer transformer decoder network with 16 attention heads and a hidden dimension of 1024. We train the transformer on WikiText with a context length of 256 and achieve a final test perplexity of 33.10.
C.1.4 Image Classification
Target Networks
Deep FCN: We design a fully-connected network consisting of 50 blocks. Each block contains a feedforward linear layer, a batch normalization, and a ReLU nonlinear activation. The intermediate feedforward linear layers contain 2048 units. This network is untrainable due to vanishing gradients since the network is very deep and due to overfitting.
Wide FCN: We design a network similar to Deep FCN but only containing 3 blocks where each feedforward linear layer contains 8192 units. This network is considered untrainable due to a saturation on the training performance.
Deep ConvNet: We design a deep convolutional network with the same architecture as ResNet-50 (convolutional layers followed by batch normalization) but remove the residual connections. This makes the network untrainable due to the vanishing gradient problem as observed in He et al. [31], causing saturation of the loss.
Guide Networks
ResNet-18/50: A deep convolutional network with 18/50 convolutional blocks and residual connections. We refer to He et al. [31].
We supervise the Deep FCN and Wide FCN with ResNet-18 and supervise the Deep ConvNet with ResNet-50.
C.2 Training
Tasks Experiment Learning Rate
Copy-Paste RNN $1\times 10^{-4}$
Transformer $1\times 10^{-4}$
Transformer $\rightarrow$ RNN $1\times 10^{-4}$
Untrained Transformer $\rightarrow$ RNN $1\times 10^{-4}$
Parity Transformer $1\times 10^{-3}$
RNN $1\times 10^{-2}$
RNN $\rightarrow$ Transformer $1\times 10^{-3}$
Untrained RNN $\rightarrow$ Transformer $1\times 10^{-3}$
Language Modeling Small RNN $1\times 10^{-4}$
Small Transformer $1\times 10^{-4}$
Small Transformer $\rightarrow$ Small RNN $1\times 10^{-4}$
Untrained Small Transformer $\rightarrow$ Small RNN $1\times 10^{-4}$
Large RNN $1\times 10^{-4}$
Large Transformer $1\times 10^{-4}$
Large Transformer $\rightarrow$ Large RNN $1\times 10^{-4}$
Untrained Large Transformer $\rightarrow$ Large RNN $1\times 10^{-4}$
Image Classification Deep FCN $1\times 10^{-4}$
Wide FCN $1\times 10^{-4}$
Deep ConvNet $1\times 10^{-3}$
ResNet-18 $\rightarrow$ Deep FCN $5\times 10^{-5}$
Untrained ResNet-18 $\rightarrow$ Deep FCN $5\times 10^{-5}$
ResNet-18 $\rightarrow$ Wide FCN $1\times 10^{-4}$
Untrained ResNet-18 $\rightarrow$ Wide FCN $1\times 10^{-4}$
ResNet-50 $\rightarrow$ Deep ConvNet $1\times 10^{-3}$
Untrained ResNet-50 $\rightarrow$ Deep ConvNet $1\times 10^{-3}$
Table 4: Learning rates for network training. For all networks, we sweep over 5 learning rate values before choosing the learning rate with the lowest validation loss for training. Our training does not use any learning rate scheduling such as a warm-up scheduler although such techniques may improve results.
In Table 4, we show the different learning rate settings we converged to in each experiment. For each experiment, we did a grid search over 5 different learning rate parameters to ensure optimal learning rate setting. We did careful tuning of all training of target networks to ensure maximum performance.
For all image classification tasks, we used the Adam optimizer [43], in-line with prior work [31]. For all sequence modeling tasks, we use AdamW [52], which has been useful in training sequence models like RNNs and Transformers [61].
The training experiments in this paper were completed across 4 H100s and 4 A100 GPUs for 3 weeks in total. GPU optimization techniques were taken such as gradient accumulation and gradient checkpointing and some language modeling experiments used mixed-precision training.
Appendix D Representational Regularization: Guidance with Noise
Figure 6: Feeding noise prevents overfitting. We introduce an additional experiment where we feed noise to our guide network rather than the same batch during each training step with guide network guidance. We sample noise from a Gaussian distribution with a mean of 0 and a standard deviation of 1. We find that despite having no information about the images in the batch, the guide network still provides an inductive bias to prevent overfitting. While this noise increase the training loss, this shows a true transfer of an inductive bias that is not driven by pure distillation of similar features.
We also aim to understand the role of the guide network as in guidance. In all experiments, we use trained and untrained guide networks and see consistent improvements for training the target network. The success of untrained networks implies that our training method is not performing distillation but instead truly transferring a prior from the guide network to the target network. To more strictly test this theory, we include an experiment where we feed noise to the guide network instead of the same batch of data fed to the target network as implied by eq.˜1.
We apply this experiment to the Deep FCN with an untrained ResNet-18 as the target network. At each training step, we pass a noisy batch which is sampled from a random Gaussian with mean of 0 and standard deviation of 1. We train for 100 epochs and report the learning curve results in fig.˜6.
This result confirms our intuition about the role of guide network: as a guide on model priors rather than a pure distillation of information. While the overall cross-entropy loss magnitudes are higher and the overall accuracy is lower when passing noise to the guide network, our results are significantly better than applying vanilla training approaches to the Deep FCN.
Appendix E Representational Similarity Loss
Figure 7: CKA representational similarity loss for sequence modeling tasks. We visualize the total CKA dissimilarity loss across all layers across training for all three sequence modeling tasks. The CKA dissimilarity loss represents the representational alignment between our guide network and target network. We can observe that for the copy-paste task and language modeling task, the target network aligns with a randomly initialized network more quickly. This could be because of special properties of RNNs.
Figure 8: CKA representational similarity loss for image classification. We visualize the total CKA dissimilarity loss across all layers across training for image classification. The CKA dissimilarity loss represents the representational alignment between our guide network and target network. We can observe that for Deep FCN and Wide FCN, the target network aligns with a randomly initialized network more quickly. This corresponds with results where randomly initialized guide networks had superior performance to trained guide networks.
We can view the representational alignment between the guide and target networks during training. This allows us to better understand how this representational alignment influences network performance. We show sequence modeling results in fig.˜7 and image classification results in fig.˜8.
We notice that across most tasks, reducing representational dissimilarity is easier with activations from randomly initialized networks rather than trained networks. This provides additional evidence of representational alignment for inductive bias transfer. We notice that for certain cases, such as Parity, the randomly initialized guide network has higher representational dissimilarity loss than the trained guide network. This is matched with the Parity result in table˜2 and fig.˜2.
However, we can also observe more inconsistent results with the Deep ConvNet where the untrained guide network has lower representational dissimilarity loss than the trained guide network, even at the end of training. One possible explanation that the inductive bias was more similar for Deep ConvNet and ResNet-50. This means that trained features are more important for better Deep ConvNet results and representational alignment with a trained network is important. This result has interesting implications for understanding the role of residual connections. Since untrained ResNet-50 is easier to align with than a trained ResNet-50, this demonstrates that residual connections influence representation spaces during training. The untrained residual connections have little influence on the inductive biases of the network or the overall representation space. This demonstrates the strength of our method as a way to interpret neural network design choices and how they influence representation and functional aspects of a network.
These results are also potentially indicative of architectural properties of RNNs and FCNs which match randomly initialized networks more quickly. For instance, one potential explanation is that RNNs have more degrees of freedom [8] and therefore, only need inductive guidance rather than trained features. Transformers may require learned features indicating that the bottleneck for transformers on the parity is not algorithmic but feature-based. Much of the future work can use these results to design better networks with more informed designs.
E.1 Layerwise Analysis
Figure 9: CKA dissimilarity decreases more in earlier network layers than later layers. When we separate the CKA dissimilarity across layers of the target network, we find that earlier layers optimize more and earlier. We take two layers from a Deep FCN during guidance with a randomly initialized ResNet-18. The early layer comes from the 15th FCN block. The later layer comes from the 43rd FCN block. We see that both layers are eventually optimized but the later layer receives less supervision and has a higher CKA at the end of training.
We provide a deeper analysis of patterns of the representational dissimilarity across different layers during guidance in fig.˜9. We find that earlier layers generally have higher CKA similarities with their corresponding layer from the guide network and later layers have lower CKA similarities. Furthermore, these later layers optimize later in the training process.
Appendix F Error Consistency
We measure error consistency $(\kappa)$ [23] between the guided target networks which indicates the error overlap between two networks based on the accuracy of the networks, i.e. do the two networks make similar class predictions? The measure first calculates the expected error overlap. Suppose $a_{1}$ is the accuracy of the first guided network and $a_{2}$ is the accuracy of the second. The expected error overlap is given by $c_{\text{exp}}=a_{1}*a_{2}+(1-a_{1})*(1-a_{2})$. Next, we measure the observed error overlap across each sample in the validation set as $c_{\text{obs}}=\text{\# of samples where both models agree}~/~\text{total trials}$. Finally, we can write $\kappa$ as:
$\kappa=\frac{c_{\text{obs}}-c_{\text{exp}}}{1-c_{\text{exp}}}$ (7)
$\kappa$ ranges from 1-1 to $1$, where $1$ is perfect agreement, 1-1 is perfect disagreement and 0 is change agreement. When $\kappa>0$, this implies that models make consistent error patterns, $\kappa<0$ implies that models make inverse error patterns, and $\kappa\approx 0$ implies independent error patterns.
F.1 Guide Network Representation Comparison
Figure 10: Earlier layers of ResNet-18 and ViT-B are more similar. We analyze the representational similarity between activations from layers ResNet-18 and ViT-B-16 via CKA. We find that earlier layers are more similar while later layers have divergent representations. We see that this manifests in distinct error consistency patterns when ResNet-18 and ViT-B are used as guide networks.
We contextualize the findings in error consistency by comparing the representations of the guide networks, in this case ResNet-18 and ViT-B-16. We apply a layer mapping between layers of ResNet-18 and ViT-B-16 and compute the representational similarity over 1000 input images. Results are shown in fig.˜10.
Our findings are useful for error consistency. If models are inheriting inductive biases from their guide network, then the models would have similar methods to process low-level image features are indicated by a stronger CKA in earlier layers between the ResNet-18 and ViT-B. This means that errors will be consistent for low level features but inconsistent for high level features collected in later layers.
Appendix G Test Accuracy across Training
Figure 11: Testing accuracy improves across guidance for sequence modeling. We visualize the testing accuracy for sequence modeling as an example to demonstrate that guidance improves accuracy across training and this improvement is significantly better across training. This allows for another interpretation of the method outside of cross-entropy loss.
We plot accuracies over training as a complement to cross-entropy loss in fig.˜11 for the sequence modeling experiments. We can use these experiments to de-couple our results from properties of cross-entropy loss that may lead to misleading improvements across training. We find that accuracies improve consistently across training, supporting the loss curve interpretation that guided training improves results.
Appendix H Basic Distillation Comparison
Figure 12: Distillation does not prevent overfitting. We compare basic distillation [33] to see if we can prevent overfitting. We use Deep FCN as our student network. We find that distillation with a trained ResNet-18 teacher network leads to a small improvement in performance but still has some patterns of overfitting. Distillation with an untrained ResNet-18 teacher network hurts performance.
To show the effectiveness of guidance, we compare it with distillation from [33]. Distillation involves transferring knowledge from a performant teacher network to a less performant student network via maximizing the alignment of the output logits. This encourages the student to have similar predictions as the teacher network. This occurs via the following loss function. Assume $\bm{Q}$} is the logits extracted from the target (student) network and the $\bm{P}$} is the logits extracted from the guide (teacher) network.
$\mathcal{L}_{\text{distill}}=\alpha*T^{2}*\text{KL}(\sigma({\bm{Q}}/T)||\sigma({\bm{P}}/T))+(1-\alpha)*\mathcal{L}_{\text{CE}}({\bm{Q}},y)$ (8)
where $y$ is the ground truth labels, $T$ is the temperature to soften the logits, and $\alpha$ is the weighting factor between the distillation loss and cross-entropy loss. In this case, KL refers to the Kullback-Liebler divergence and $\sigma$ corresponds with the softmax function. In practice, we set $\alpha$ to 0.5 and $T$ to 2. We continue to track the full cross-entropy loss across training as well.
We show accuracy-based results in fig.˜5 and loss curve results in fig.˜12. Distillation with a trained network can improve performance but much less than guidance. Distillation with an untrained network reduces performance on average, although not by a significant amount.
H.1 Error Consistency
Figure 13: Distillation lowers error consistency. In general, we find that distillation results in error consistency patterns that are less consistent than what is reported with guidance.
We show error consistency performance over distilled networks rather than guided networks in fig.˜13.
Appendix I Guided Network Analysis and Interpretation
The results from guidance open many questions in order to explain why untrained guide networks can be better at improving target network performance. We provide an intuitive explanation as well as some geometric analysis of guided networks to see if there is a stronger interpretation.
I.1 Interpreting Guidance
We believe several prior works support our findings and interpretations in this paper. We cover them here. First, [38] is a recent paper that considers task-aware representational alignment. Their theory provides a generalization bound via kernel alignment. They show that when a “stitcher” maps representations or a source network to a target output, the excess risk of the stitched model is upper-bounded by the CKA alignment between them. This provides a learning-theoretic guarantee that the CKA term in guidance reduces the hypothesis class possibilities seen by an optimizer. Overfitting or underfitting becomes harder.
Shan and Bordelon [70] investigates how a network’s neural tangent kernel (NTK) aligns with a target output during training. The paper shows that NTK alignment accelerates convergence and lowers generalization error in deep linear networks. This aligned kernel condition is inserted by hand in guidance. Similarly, [3] uses Rademacher complexity tools to show that alignment of tangent-kernel features onto a small set of task-relevant directions compresses the effective model class. This formalizes the notion of guidance as an automatic regularizer, where task directions are replaced by the guide network settings. Finally, [37] demonstrates that after training, the top singular vectors of a network’s hidden activations align with the task target vectors. This empirically supports the layerwise CKA choice in guidance. We believe that CKA bounds the risk or complexity in terms of kernel alignment. The NTK and Rademacher analyses show that alignment shrinks the effective hypothesis space and improves conditioning. This aligns with findings based on singular vectors. We could sharpen this theory by changing the alignment used in guidance e.g. moving from aligning on kernels to aligning on singular vectors or eigenvectors instead. A full PAC-style proof specialized to guidance has not been shown in our paper but we leave this to future work.
I.2 Geometric Analysis via Intrinsic Dimensionality
Figure 14: Guidance preserves intrinsic dimensionality, avoiding over-regularization. We measure the PCA-based intrinsic dimensionality of representations from each layer of both a guided and unguided Deep FCN at initialization i.e. Random FCN, 300 steps of training, and the end of training. We find that guidance with an untrained guide network better preserves intrinsic dimensionality in comparison to base training or using a trained guide. All networks collapse to the same intrinsic dimensions. This establishes that guidance does change the dynamics of training based on geometric features of the target network.
We aim to understand how guidance with a randomly initialized guide network differs from a trained guide network. To do so, we compare the representation space of a target network guided by a trained guide network and a randomly initialized guide network using intrinsic dimensionality.
Intrinsic dimensionality (ID) refers to the minimum number of dimensions required to capture the structure or variability in the input data. It represents the true complexity of the data manifold, ignoring noise or redundant dimensions. Previous work has found that neural networks have low ID, capturing data in low-dimensional manifolds [48, 60]. Following [22], for a given threshold $\beta$, the intrinsic dimension is the $d\in\mathbb{N}$ such that the ratio of explained variance for $d$ dimensions of a $N$ dimensional PCA is above $\beta$:
$\frac{\sum_{i=1}^{d}var(y_{i})}{\sum_{j=1}^{N}var(y_{j})}>\beta$ (9)
First, we measure the ID of both unguided and guided Deep FCNs across all layers at different points of training; see fig.˜14. Crucially, at 300 steps of training, we notice that guidance with an untrained guide network preserves the initial ID found in the randomly initialized Deep FCN representations. The trained guide network and base training achieve low ID values, which is consistent with findings in prior work [11]. At the end of training, all networks reach the same intrinsic dimension. This finding indicates that the dynamics of training, as shown by ID, change with or without guidance. The Deep FCN without guidance achieves a low ID too early in training, and this is likely similar for the guidance with a trained guide network.
One interpretation of these results is that ID has an effect on overfitting in the Deep FCN. When ID is too low, the Deep FCN overfits. Therefore, we can discover a new regularization scheme for training the Deep FCN based on ID. In this scheme we introduce a new loss function, which designs a differentiable version of PCA-ID based on a specific ID threshold and forces the ID of the representations to be above a particular ID. In particular, given the activations from a specific layer of the target network, ${\bm{A}}^{T}_{i}$, a variance threshold $\tau$, and a target ID $t$, we first find the SVD of the activations,
${\bm{A}}^{T}_{i}={\bm{U}}{\bm{\Sigma}}{\bm{V}}^{T}$ (10)
We extract the eigenvalues using the singular values, $\lambda_{j}=\Sigma_{jj}^{2}$ and find the total variance using the eigenvalues.
$T=\sum_{j=1}^{r}\lambda_{j}+\epsilon$ (11)
where $r$ is the total number of nonzero singular values and $\epsilon$ is a small constant for numerical stability. Afterwards, we find the explained ratios and cumulative sums of the eigenvalues:
$p_{j}=\frac{\lambda_{j}}{T}~,~c_{k}=\sum_{j=1}^{k}p_{j}$ (12)
We use the cumulative sum to find the loss for being below the target ID in eq.˜13.
$\ell_{\text{below}}=\sum_{j=1}^{t-1}\sigma(\beta(c_{j}-\tau))$ (13)
where $\sigma$ is the sigmoid function and $\beta$ controls the sharpness of the loss. Similarly, eq.˜14 gives the loss for being too far above the target ID.
$\ell_{\text{above}}=\sigma(\beta(\tau-c_{t}))$ (14)
Figure 15: Controlling intrinsic dimensionality of Deep FCN representations improves image classification performance. We introduce a novel loss function to regularize the ID of Deep FCN representations in each linear layer of the network during training. We find that this leads to improved validation accuracy, similar to guidance. This shows that guidance correlates with geometric modifications to target networks and can find new regularizations.
Our total loss is given by $\mathcal{L}=\ell_{\text{above}}+\ell_{\text{below}}$. Using this loss to control ID, we tune each layer in our Deep FCN and see a final validation loss given in fig.˜15. We find that using our new loss function to control the ID of every linear layer of our Deep FCN leads to improved validation performance and an accuracy of 14.65%. This aligns with guidance, indicating that guidance may be controlling the ID of the target network. Furthermore, we have used guidance to find a new regularization scheme based on intrinsic dimensionality.
I.3 Linear Decoding
Figure 16: ImageNet classes are barely decodable from a randomly initialized ResNet-18. In order to assess the performance of our randomly initialized networks, we design a linear decoder to decode ImageNet classes from all layers of the network. Chance accuracy is 0.1% (1/1000) on the graph above. We find that, while we can decode ImageNet classes with an accuracy above chance, the accuracy is still very low.
We assess whether object classes are decodable from internal representations of a randomly initialized ResNet-18. A potential explanation for improvements in the target network with an untrained guide is the ability to linearly decode the ImageNet classes with significant accuracy at certain layers of the untrained guide network.
We train a linear decoder with 4000 ImageNet images from the train set and test on 1000 images from the validation set for each layer. We show results in fig.˜16. We find that the linear decoder never achieves any accuracy greater than 0.8% for any of the layers. Furthermore, later layers contain little information that is useful to linearly decode ImageNet classes. This means that linearly decodable information isn’t present in the guide network and this aspect isn’t driving improvements in target networks. We note that this matches findings in Amid et al. [2], which reports that linear decodability from a ResNet-18 achieves 3.4% top-1 accuracy. The increase in performance is likely due to using a much larger dataset.
Appendix J Guidance with RSA and Ridge Regression
J.1 Representational Similarity Analysis
We use the RSA formulation as described in Kriegeskorte et al. [46]. Specifically, RSA constructs representational dissimilarity matrices (RDMs) for two sets of representations and compares them using an outer similarity function.
Given two sets of representations, ${\bm{R}}\in\mathbb{R}^{b\times d_{1}}$ and ${\bm{R}}^{\prime}\in\mathbb{R}^{b\times d_{2}}$, we first calculate RDMs for each set of representations using a distance function $d$. Formally, we define ${\bm{D}}\in\mathbb{R}^{b\times b}$ as
${\bm{D}}_{i,j}:=s({\bm{R}}_{i},{\bm{R}}_{j})$ (15)
Each row ${\bm{D}}_{i}$ corresponds to the distance between the representations of input $i$ and the representations of all inputs including itself. This is done per-batch, meaning that RSA is sensitive to batch size.
Given two RDMs $\bm{D}$} and $\bm{D}$$^{\prime}$ constructed from sets of representations $\bm{R}$} and $\bm{R}$$^{\prime}$ respectively, we vectorize the RDM matrices using a function $v$ (since the RDMs are symmetric, we only need to compare the lower triangles), and compute the similarity between the two vectorized RDMs using a similarity function $s$.
$\mathcal{M}({\bm{R}},{\bm{R}}^{\prime})=s(v({\bm{D}}),v({\bm{D}}^{\prime}))$ (16)
As with CKA, we use the complement of the similarity to construct $\bar{\mathcal{M}}$. In practice, we define $d$ to be the cosine distance between every pair of inputs and $s$ to be the pearson correlation between the RDMs as done in previous work [14].
We apply guidance with RSA to Deep FCN as our target network and ResNet-18 as our guide network. Similar to our CKA results, we train for 100 epochs with a batch size of 256, as RSA is sensitive to the number of samples when comparing sets of representations.
J.2 Ridge Regression
We used similar ridge regression formulation as [14, 71] without cross-validation.
Given two sets of representations, ${\bm{R}}\in\mathbb{R}^{b\times d_{1}}$ and ${\bm{R}}^{\prime}\in\mathbb{R}^{b\times d_{2}}$, we first apply a sparse random projection on the representations. Since the dimensionality of the representations is prohibitively large, the projection makes the ridge regression feasible to compute. We refer to the resulting representations as $\bm{P}$} and $\bm{P}$$^{\prime}$ which correspond to the projected representations $\bm{R}$} and $\bm{R}$$^{\prime}$ respectively. $\bm{P}$} and $\bm{P}$} have dimension $d$, where $d$ is fixed using the Johnson-Lindenstrauss lemma [40].
Afterwards, we mean-center the representations and apply ridge regressions using the original least-squared solution as follows. Our goal is to predict the representations $\bm{P}$$^{\prime}$ using regressors over $\bm{P}$}. We first split our representations into a training set i.e. ${\bm{P}}_{\text{train}},{\bm{P}}^{\prime}_{\text{train}}$ and testing set ${\bm{P}}_{\text{test}},{\bm{P}}^{\prime}_{\text{test}}$ where the training set contain half the representations and the testing set contains the other half. We first a set of regressors $\hat{\beta}$ as follows:
$\hat{\beta}=(({\bm{P}}_{\text{train}})^{T}{\bm{P}}_{\text{train}}+\lambda{\bm{I}}_{d})^{-1}({\bm{P}}_{\text{train}})^{T}{\bm{P}}^{\prime}_{\text{train}}$ (17)
where $\lambda$ is the ridge penalty, which is a hyperparameter. The coefficients $\hat{\beta}$ are then used to predict the held out data where:
$\begin{gathered}\hat{{\bm{P}}^{\prime}_{\text{test}}}={\bm{P}}_{\text{test}}\beta\end{gathered}$ (18)
We measure the cosine similarity between the predicted representations $\hat{{\bm{P}}^{\prime}_{\text{test}}}$ and actual representations ${\bm{P}}_{\text{test}}$.
$\mathcal{M}({\bm{R}},{\bm{R}}^{\prime})=\text{cosine}(\hat{{\bm{P}}^{\prime}_{\text{test}}},{\bm{P}}^{\prime}_{\text{test}})$ (19)
We apply guidance with RSA to Deep FCN as our target network and ResNet-18 as our guide network. Similar to our CKA results, we train for 100 epochs with a batch size of 256, as ridge regression is sensitive to the number of samples when comparing sets of representations. We manually tune the $\lambda$ hyperparameter, finding that $\lambda=10.0$ is optimal for the trained guide network and $\lambda=100.0$ is optimal for the untrained guide network.
J.3 Results
Figure 17: Guidance with RSA as the representational similarity metric maintains similar performance to CKA. We include a further experiment where we change the metric for representational alignment from CKA to RSA during guidance training. We apply this to the Deep FCN with ResNet-18 as a guide network. We see that, like CKA, RSA alignment also allows for transferring the prior from ResNet-18. However, unlike CKA, the untrained guide network only does marginally better than the trained network, potentially indicating the RSA is better at transferring trained features.
Figure 18: Guidance with ridge regression as the representational similarity metric improves performance over CKA. We change the metric for representational alignment from CKA to ridge regression during guidance training. We apply this to the Deep FCN with ResNet-18 as a guide network. We see that, like CKA, ridge regression alignment also allows for transferring the prior from ResNet-18. We find that this improves over CKA significantly
RSA: We see results over the training, validation, and representational dissimilarity loss in fig.˜17. The Deep FCN guided by a trained ResNet-18 achieves an accuracy of 11.02% and the Deep FCN guided by a randomly initialized ResNet-18 achieves an accuracy of 11.74%.
We can first observe that guided training improves over base training as noted in fig.˜2 and table˜3. This demonstrates the generality of our approach to other metrics. As long as a representational similarity metric is differentiable, we can optimize the metric for alignment between two networks as a method to transfer the prior of one network to another.
We can also observe some minute differences between the results with CKA. Most notably, the trained guide network has similar performance to the untrained guide network. This is likely because less information about trained features are present in the RSA metric. RSA measures relative distance between input instances and imposes a constraint of placing these into relative distances. It could be possible that fewer degrees of freedom are useful for aligning target network with trained guides.
Ridge: We see results over the training, validation and representational dissimilarity loss in fig.˜18. The Deep FCN guided by a trained ResNet-18 achieves an accuracy of 9.46% and the Deep FCN guided by a randomly initialized ResNet-18 achieves an accuracy of 15.69%. Similar to RSA and CKA, we can see the guided training improves over base training.
Similar to CKA, we observe that randomly initialized guide networks outperform trained guide networks. Furthermore, performance with ridge regression is better than with CKA. This finding is intuitive. Ridge regression generally has more degrees of freedom than other similarity metrics because of fewer invariances imposed on the metric. This means that the solution search space is larger, leading to better results. We believe this provides a promising path forward for making target networks have better performance.
Furthermore, ridge regression has other desirable properties such as potential explainability via probing predicted representations to measure similarity. We can use probing analyses on predicted representations to see what information the target has inherited from the guide network. This opens up many avenues for studying guidance in the future.
Appendix K Ablation Experiments
Experiment CIFAR-10 Test Accuracy $(\uparrow)$
Deep FCN 60.58
All layer Guidance 70.15
Last layer Guidance 67.18
Last two layers Guidance 68.03
Last five layers Guidance 67.50
Last ten layers 69.31
First ten layers 79.22
First five layers 79.58
First two layers Guidance 73.34
First layer Guidance 65.11
Multiple Guide Layers 68.14
Table 5: Guiding earlier layers of deep networks leads to better results. We apply an ablation experiment to identify which layers lead to stronger improvement when guided. We use a Deep FCN, guided by a randomly initialized ResNet-18 on CIFAR-10. We find that guiding earlier layers leads to strong improvement, even over guiding all layers. Guiding any layer leads to an improvement of performance.
We analyze the current design of our layer mapping for guidance by experimenting with the number of layers used in guidance and whether more complex mappings exist like mapping several layers of the guide network to a single target network.
We run layer-wise ablation experiments the Deep FCN guided by an untrained ResNet-18 over CIFAR-10. Similarly, we experiment with RNNs guided by untrained transformers over the copy-paste task.
Experiment Copy-Paste Accuracy $(\uparrow)$
RNN 14.35
All layer Guidance 42.56
Last layer Guidance 38.19
Last three layers Guidance 42.33
Last two layers Guidance 41.55
First layer Guidance 27.59
First two layers Guidance 28.15
First three layers Guidance 33.93
One guide layer 36.11
Table 6: Guidance of later layers improves RNN performance. We apply an ablation experiment over RNNs trained for copy-paste to see whether guiding certain layers lead to improved performance. We find that guiding later layers leads to stronger performance overall. Furthermore, RNN layers are guided by several guide network layers in the transformer such as the linear layer and layer-normalization in the transformer decoder. Including both of these leads to better results.
In table˜5, we first show the effect of guiding over a subset of layers in the Deep FCN, evaluated over CIFAR-10. We find that earlier layers are much more impactful for a deep network. Intuitively, this could be due to guidance providing aiding with the credit assignment problem in deep networks: gradients don’t propagate properly to earlier layers. However, table˜6 shows that later layers in the RNN are more useful to apply guidance to when improving copy-paste performance. In general, we find that guiding any layer leads to improvements in results generally, showing the general applicability.
Furthermore, in table˜5 and table˜6, we consider new methods to map guide network layers to target network layers. When guiding a Deep FCN with ResNet-18, we only apply a 1-1 layer mapping for supervision i.e. each Deep FCN layer is guided by only one ResNet-18 layer. From table˜5, introducing multiple sources of supervision from guide network layers by allowing a one-to-many mapping decreases performance. However, with the RNN, we guide with representations from the linear layer and layer normalization in the transformer decoder. One could consider that linear layers are redundant with layer normalization for guidance, so we remove its representations as a potential supervisory target. We find that this hurts performance, showing that RNNs benefit from multiple levels of supervision from its guide.
Understanding the dynamics of guidance supervision is interesting and could allow for understanding training dynamics of neural networks or allow us to form cross-architectural relationships.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We introduce guidance as a tool for taking neural networks that are considered ill-suited for specific tasks and making them trainable by aligning their activations with a trainable network like a transformer. We specifically design guidance for a suite of networks like fully-connected networks and recurrent neural networks and show systematic improvements. We state these findings in the abstract and introduction and use these to make claims about deeper understanding neural networks and potential for designing new architectures.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper addresses all limitations at the end of the introduction as well as in the appendix. We take care to make point out assumptions of our results but also establishes how we identified failures in architectures.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: This paper is empirical by nature. We do not rely on proofs. We do formulate loss functions with particular properties. These loss functions borrow from prior literature such as prior work that use neural distance functions so there are few theoretical derivations necessary.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We upload code associated with all experiments in this paper. Furthermore, we also describe all architectures in detail, include all hyperparameters used in the paper such as batch size, learning rate, and optimizer, and cover number of training steps. See our appendices or section˜4 where we cover our training setting. We run all experiments with open-source datasets that are widely available or can be easily generated.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: See previous answer. We upload code and also only work with open datasets or generated datasets. All data is available at this time.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We either use datasets with standard train/test splits or describe in detail how we generate the training/validation/testing data for generated datasets like parity or copy-paste.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We train with multiple seeds for all runs of base training and guidance and average over seeds to find the average training performance and standard error. We make all plots with this average performance and standard error for comparison of statistical significance.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We describe the compute (4 H100s) used for this experiment as well as the amount of time needed to get the experiments to work with our code set up.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We conform with all aspects of the Code of Ethics; our data does not involve human subjects or have privacy concerns. We also do not believe this work will have harmful impacts and believe this work will have positive impacts on neural architecture design.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Our conclusion is dedicated to potential impacts of our paper, with societal impacts as well.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Our paper does not require any safeguards since we are not in a high risk category.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We provide proper citation for the creators of ImageNet and Wikitext-103. We provide proper citation for the inspiration to use other tasks such as parity and copy-paste.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: We do not introduce/release new assets.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not crowdsource or conduct research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not use any crowdsourcing nor research with any human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: We do not use LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.